# Prompt-Counterfactual Explanations for Generative AI System Behavior

### v1

@@ A.1 Detection of political leaning
Table 9: Counterfactual explanations for the detection of political leaning for two models (LLaMA and OLMo). The score measures how often the output is classified as right leaning across 10 runs. In the explanations, every row presents a different explanation that brings the score (shown after the colon) below the threshold when those words are masked from the input prompt. We use $5/10$ as the threshold. The “/” indicates that the score did not exceed the threshold, so no explanations were retrieved.
@@
ABC¡¯s ¡®Historic¡¯ Diversity Hire Out Of Top Spot Amid Rumors Her Incompetence Shocked Higher Ups 7/10 Top: 0/10 Amid: 0/10 Her: 0/10 Incompetence: 0/10 Shocked: 0/10 0/10 /
Columbia cancels university-wide graduation ceremony after weeks of anti-Israel protests 10/10 protests: 5/10 Columbia, cancels: 3/10 2/10 /
Boeing 787 employees falsified inspection records; FAA opens probe 7/10 787: 0/10 inspection: 0/10 opens: 0/10 probe: 0/10 falsified: 4/10 0/10 /
”We are screwed”: 2nd House Republican joins Mike Johnson ouster effort 7/10 joins: 0/10 are: 4/10 ouster: 4/10 Mike, Johnson: 4/10 5/10 /
Arrested US soldier awaiting hearing in Russia on theft charges 9/10 soldier: 1/10 charges: 2/10 awaiting: 3/10 theft, hearing: 3/10 8/10 on: 0/10 theft: 0/10 charges: 0/10 in: 1/10 US: 2/10
Biden delivers major speech on antisemitism at Holocaust remembrance ceremony 10/10 Biden: 4/10 5/10 /
NPR editor¡¯s tell-all confirms what we already knew about the media 3/10 / 1/10 /
Biden says he¡¯s ¡®happy¡¯ to debate Trump 7/10 he¡¯s: 3/10 ¡®happy¡¯: 4/10 to: 5/10 5/10 /
US intel signals Putin not directly to blame for Navalny¡¯s untimely death: Reports 6/10 not: 2/10 for: 2/10 blame: 3/10 directly: 4/10 Navalny¡¯s: 4/10 6/10 not: 0/10 directly: 0/10 Navalny¡¯s: 0/10 to: 2/10 blame: 3/10
Kremlin Disputes Report Putin Didn¡¯t Order Navalny¡¯s Death 10/10 Death: 1/10 Report: 5/10 Putin: 5/10 6/10 Putin: 1/10 Navalny¡¯s: 1/10 Disputes: 2/10 Death: 2/10
Read: Trump hush money gag order ruling 10/10 ruling: 2/10 Trump: 3/10 money: 5/10 3/10 /
House Democrats announce they would save Speaker Mike Johnson if Marjorie Taylor Greene triggers her effort to oust him 6/10 announce: 1/10 if: 3/10 Democrats, Speaker: 0/10 7/10 announce: 0/10 Johnson: 1/10 Taylor: 2/10 Greene: 2/10 her: 2/10
Supreme Court Appears Skeptical of Using Obstruction Law to Charge Jan. 6 Rioters 7/10 Rioters: 1/10 5/10 /
Georgia parliament descends into chaos as lawmakers throw punches over Putin-style ¡®foreign agent¡¯ bill 5/10 / 2/10 /
¡®Journalistic Rape¡¯: Catherine Herridge Accuses CBS News of Crossing a ¡®Red Line¡¯ During House Hearing 8/10 Accuses: 0/10 During: 0/10 Rape¡¯:: 1/10 a: 2/10 Crossing: 3/10 3/10 /
It¡¯s No Secret Our Tax System Punishes Low-Income People. It Doesn¡¯t Have To. 2/10 / 10/10 To.: 0/10 Our: 1/10 Doesn¡¯t: 1/10 System: 2/10 Punishes: 2/10
US moving ’assets’ to region to deter Iran from retaliatory attack on Israel, avoid wider conflict, officials say 6/10 say: 1/10 moving: 4/10 deter: 4/10 on: 5/10 ’assets’, Israel,: 1/10 6/10 ’assets’: 0/10 to: 0/10 deter: 0/10 conflict,: 0/10 officials: 0/10
House Democrats will block motion to remove Speaker Johnson 2/10 / 7/10 House: 1/10 to: 4/10 Democrats: 5/10
@@ A.3 Sentiment analysis
Table 11: Counterfactual explanations for sentiment analysis for two models (LLaMA and OLMo). The score measures the average negative sentiment of the output generated by that prompt across 10 runs. In the explanations, every row presents a different explanation that brings the score (shown after the colon) across the threshold when those words are masked from the input prompt. We use 0.00 as the threshold. The “/” indicates that the score did not exceed the threshold, so no explanations were retrieved. We only provide the first two rows of the results.

### v2

@@ A.1 Detection of political leaning
Table 9: Counterfactual explanations for the detection of political leaning for two models (LLaMA and OLMo). The score measures how often the output is classified as right leaning across 10 runs. In the explanations, every row presents a different explanation that brings the score (shown after the colon) below the threshold when those words are masked from the input prompt. We use $5/10$ as the threshold. The “/” indicates that the score did not exceed the threshold, so no explanations were retrieved. We only provide a subset of the results due to space constraints.
@@ A.3 Sentiment analysis
Table 11: Counterfactual explanations for sentiment analysis for two models (LLaMA and OLMo). The score measures the average negative sentiment of the output generated by that prompt across 10 runs. In the explanations, every row presents a different explanation that brings the score (shown after the colon) across the threshold when those words are masked from the input prompt. We use 0.00 as the threshold. The “/” indicates that the score did not exceed the threshold, so no explanations were retrieved. We only provide the first two rows of the results due to the journal’s space constraints on submitted manuscripts.